OpenStudy (anonymous):

1. Roman mixes 12 liters of 8% acid solution with a 20% acid solution, which results in a 16% acid solution. Find the number of liters of 20% acid solution in the new mixture. Construct a table to organize the information, and show all of your work to solve the problem. Really need help with this ASAP its for a semester test... Please explain in details please and don't just say the answer...

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

guess we need a variable here right? lets call \(x\) the number of liters of 20% solution

OpenStudy (anonymous):

Hai!

OpenStudy (misty1212):

then the amount of acid will be \[.08\times 12+.2\times x\] and the total amount is \(12+x\) which is \(16\%\) acid so \[.08\times 12+.2\times x=.16\times (12+x)\] and you can solve this for \(x\)

OpenStudy (anonymous):

OK, I need this answer to be step by step like First: we do etc,etc,etc, with the variable. Thank you so much for helping me! :D

OpenStudy (anonymous):

Why is it copying what we're saying and reposting them?

OpenStudy (anonymous):

Are you still there?